OpenStudy (anonymous):

√35 OVER √38 Do i just multiply both sides by √38 then simplify?

OpenStudy (anonymous):

i was told you cannot have a square root on the denominator and to get rid of it, you multiply both sides by the denominator.

OpenStudy (asnaseer):

@maherman97 you are correct

OpenStudy (anonymous):

so it would be √1330 OVER √1444

OpenStudy (asnaseer):

well, you can simplify the denominator here

OpenStudy (asnaseer):

because remember you got to it by doing this:\[\frac{\sqrt{35}}{\sqrt{38}}=\frac{\sqrt{35}}{\sqrt{38}}\times\frac{\sqrt{38}}{\sqrt{38}}=\frac{\sqrt{35*38}}{\sqrt{38*38}}\]

OpenStudy (asnaseer):

so what is \(\sqrt{38*38}=\sqrt{38^2}=?\)

OpenStudy (asnaseer):

@micahwood50 that is not how radicals are simplified. @maherman97 has the correct approach for this type of problem

OpenStudy (asnaseer):

@maherman97 what is \(\sqrt{38*38}=\sqrt{38^2}=?\)

OpenStudy (asnaseer):

i.e. what is the square root of something squared?

OpenStudy (asnaseer):

more generally, what is \(\sqrt{x^2}=?\)

OpenStudy (anonymous):

SOO SORRY! its 38! so i would have √1330 OVER 38???

OpenStudy (asnaseer):

yes - that is the correct answer. :)

OpenStudy (anonymous):

and i would simplify?

OpenStudy (asnaseer):

that is now in its simplest formas \(\sqrt{1330}\) cannot be simplified any further

OpenStudy (anonymous):

well what about like 2 and 665

OpenStudy (anonymous):

and so on

OpenStudy (asnaseer):

\[1330=2\times5\times7\times19\]all of which are prime number, so the square root cannot be simplified any further

OpenStudy (anonymous):

ohhh OK THANKS SOO MUCH!

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

i have another!

OpenStudy (asnaseer):

<-- please post each question separately on the left.